If the price of a good decreases from $7 to $6 and it's quantity demanded increased from 18 to 20 , what is the price elasticity of the good


1
Expert's answer
2021-03-08T07:21:30-0500

We can find the price elasticity of the good as follows:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{(Q_2+Q_1)0.5}}{\dfrac{P_2-P_1}{(P_2+P_1)0.5}},""E_d=\dfrac{\dfrac{20-18}{(20+18)\cdot0.5}}{\dfrac{\$6-\$7}{(\$6+\$7)\cdot0.5}}=-0.68."

The price elasticity of the good is 0.68 which means that demand is inelastic.